what is the equation for the line that pass through the point (-1,4), and is perpendicular to the line with the equation x=2?

Answer by mananth(16949)

x=2 is a line parallel to the y axis
slope =0
the perpendicular line passes through (-1,4)
y= mx+b
4=0+b
b=4 the y intercept
so the equation will be y =mx+4
m=0
y=4